OpenStudy (anonymous):

A simple looking tough question pertaining to number of digits. Question below

OpenStudy (anonymous):

http://mypat.in/PaperImages/3/425/Q_14_014.gif

OpenStudy (anonymous):

5^30=(10/2)^30=p(say) now log p =log 10^30 - log 2^30 = 30-30*.3010=20.97 so 5^30 lies between 10^20 and 10^21.. so it has 21 digits

OpenStudy (anonymous):

Thanks @1729antony , good work brother. Dunno where I left my brain today :P

OpenStudy (anonymous):

:)